OpenStudy (anonymous):

If f(x) = x + 5, find f(−9). 14 4 −14 −4

jimthompson5910 (jim_thompson5910):

replace every copy of x with -9, then evaluate

OpenStudy (luigi0210):

plug in -9: f(-9)=(-9)+5

OpenStudy (anonymous):

Please help with these questions. I will try to answer as many as i can but am stuck on this one.

OpenStudy (anonymous):

Thanks. -4

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to y = −3x − 1 and that pass through the point (3, 1). parallel: y = −3x + 6 perpendicular: y = one over threex + two and two thirds parallel: y = −3x + 10 perpendicular: y = one over threex parallel: y = negative one over threex perpendicular: y = −3x parallel: y = one over threex + 1 perpendicular: y = 3x − 1

OpenStudy (anonymous):

I think that it is the top one.

jimthompson5910 (jim_thompson5910):

you need to make sure the point (3,1) is on both lines

jimthompson5910 (jim_thompson5910):

y = -3x + 6 1 = -3*3 + 6 ... plug in (3,1) 1 = -9 + 6 1 = -3 ... FALSE, so (3,1) is NOT on the line y = -3x + 6

OpenStudy (anonymous):

Ok. I got it.